write an equation of the line containing the given point and perpendicular to the given line. Type your answer in the form y=mx+b simplify your answer.
(9,-9);5x+2y=3 I do not understand how to do any of these problems, please help before I fail.

Answer by Alan3354(69443)

write an equation of the line containing the given point and perpendicular to the given line. Type your answer in the form y=mx+b simplify your answer.
(9,-9);5x+2y=3
---------------
Find the slope of the line:
5x+2y=3
Find the slope, m, by putting it in the slope-intercept form y = mx + b.
That means solve for y.
5x+2y=3
2y = -5x + 3
y = (-5/2)x + 3/2
The slope is -5/2
-------------------
The slope of lines perpendicular is the negative inverse, +2/5.
Now the problem is to find the equation of a line thru (9,-9) with a slope of 2/5.
Use y = mx+b and the point to find b:
-9 = (2/5)*9 + b
-9 = (18/5) + b
b = -9 - 18/5
b = -63/5
----------
y = (2/5)x - 63/5
That's it, in slope-intercept form.